OpenStudy (anonymous):

5 - 2/7 (show steps please & thank you!)

Directrix (directrix):

5 - 2/7 = 35/7 - 2/ 7 = 33 / 7 = 4 5/7

OpenStudy (anonymous):

Thank you @Directrix! :)

OpenStudy (amistre64):

5 -1/7 - 1/7 5-1/7 = 4 6/7 - 1/7 = 4 5/7